OpenStudy (anonymous):

Roman weight 25 kg. His Father weight 5kg lessthan 3 times roman's weight. Answer pls.

OpenStudy (jessiegonzales):

3 times romans weight is 75kg so his father is 3 less s 72kg?

OpenStudy (anonymous):

Can you give the equation ? @jessiegonzales

OpenStudy (anonymous):

the equation would be (25 x 3) - 5 Sooo... = (25 * 3) - 5 = 75 - 5 = 70 So his father's weight is 70kg! :D